OpenStudy (anonymous):

any one good at graphing quadratic functions if so respond and ill post the problems.

OpenStudy (campbell_st):

what do you need to know...?

OpenStudy (anonymous):

how to graph these: 1. f(x)=-2x 2. f)x)=x^2-4x+4 3. f(x)=x^2-6x+8

OpenStudy (campbell_st):

ok... 1. is a straight line f(x) = -2x use a table of values x: 1, 2, 3 y: -2 -4 -6 so the ordered pairs (1, -2), (2, -4) and (3, -6) could be plotted and a line drawn through them . 2. Can you factorise the quadratic...?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

what is the factorised form?

OpenStudy (anonymous):

(x-2)(x-2)

OpenStudy (campbell_st):

yep \[y = (x - 2)^2\] so the graph has a root( or contacts the x-axis) when \[(x - 2)^2 = 0\] so a root exists at x = 2 the curve has a y-intercept when x = 0 so the y- intercept is y = 4 and the parabola is concave up. The general form is \[y=ax^2 + bx + c\] in your question a > 0 so its concave up...  the curve should touch the x-axis at x = 2